OpenStudy (anonymous):

How do you integrate (sin2pix)^2 with the boundaries being 1/4 on top and 0 at the bottom?

OpenStudy (lgbasallote):

\[\huge \int_0^{1/4} [\sin (2\pi x)]^2dx\] is that the question?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

interesting...no cosines huh

OpenStudy (anonymous):

how can you write it like that? Is it a website?

OpenStudy (lgbasallote):

nope...it's a latex code...you can do some basic stuffs by using the "Equation button" below the reply box

OpenStudy (anonymous):

oh sorry, there's no 2 in the 2pix

OpenStudy (lgbasallote):

so pi x?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

you have learned integration by parts already right?

OpenStudy (anonymous):

hm.. I'm not sure what you mean - what's integration by parts?

OpenStudy (lgbasallote):

\[\huge uv - \int vdu\] or something like that...no?

OpenStudy (anonymous):

is it when you use u to replace something? like dy/du x du/dx?

OpenStudy (anonymous):

Use this \[\sin^2 \theta=\frac{1-\cos 2\theta}{2}\]

OpenStudy (anonymous):

yup I used that, but I got stuck

OpenStudy (lgbasallote):

let \(\theta\) = \(\pi x\) first

OpenStudy (anonymous):

\[\int\limits_{1/4}^{0}(1-\cos2pix)/2 dx\]

OpenStudy (anonymous):

opps I typed it wrong. Hmm.. ok let theta = \[pix\]

OpenStudy (anonymous):

it would be simpler if u let \(\theta=2\pi x\) so \(\theta\) goes from 0 to \(\pi/2\) since \(x\) goes 0 to \(1/4\)

OpenStudy (anonymous):

Ok, is \[\pi/8 - 1/4\sin(\pi/2)\] right? I'll try using \[2 \pi x\] now

OpenStudy (anonymous):

@mukushla I tried using \[2 \pi x \] but I ende dup with 1 - 0.5 ((4sin\[\pi /2 \]/pi - sin 0/0)

OpenStudy (anonymous):

Sorry that it's so messy

OpenStudy (australopithecus):

u = 2x du/2pi = dx \[\frac{1}{2\pi}\int\limits_{0}^{1/4}\sin^{2}(u)du = \frac{1}{2\pi}\int\limits_{0}^{1/4}\frac{1-\cos(2u)}{2}\] now you can easily solve this integral

OpenStudy (australopithecus):

\[\frac{1}{2\pi}\int\limits_{0}^{1/4}\frac{1}{2}du - \int\limits_{0}^{1/4}\frac{\cos(2u)du}{2}\]

OpenStudy (australopithecus):

note there should be brackets around it but yeah easy to solve now :)

OpenStudy (anonymous):

oh that makes sense now. I can solve it now! I'll tell you when I'm done!

OpenStudy (australopithecus):

yeah there should be brackets around it both integrals should have 1/2pi multiplied into them

OpenStudy (anonymous):

ok, I added the brackets and I got 1/16 - 1/8pi is it right?

OpenStudy (australopithecus):

\[\frac{1}{2\pi}\int\limits_{0}^{1/4}\frac{1}{2}du - \frac{1}{2\pi}\int\limits_{0}^{1/4}\frac{\cos(2u)du}{2}\]

OpenStudy (anonymous):

oh.. thanks. My answer is really off. Um, do I have to change the boundaries if I'm using "u"?

OpenStudy (australopithecus):

well you can but I usually prefer to just resub for u after integrating

OpenStudy (australopithecus):

so replace u with 2pix

OpenStudy (australopithecus):

after integrating

OpenStudy (australopithecus):

you will also want to do another substitution on cos(2u)/2 before integrating

OpenStudy (anonymous):

ok, I'm working on it = )

OpenStudy (australopithecus):

Here this will show you how to deal with integrals such as these

OpenStudy (australopithecus):

this is the source covers most everything done in calculus 1 at my university

OpenStudy (anonymous):

wow thanks! It summarises everything! hmm.. ok I've got 1/8 - 1/4pi now but it still doesn't equal to 1/8

OpenStudy (australopithecus):

can you scan or take a picture of your work?

OpenStudy (anonymous):

ok, I'll try write it up

OpenStudy (anonymous):